In a school of 1620 pupils the ratio of boys to girls is 4:5
How many girls in the school?
Teaching a 10 year old so easy as possible solution
Thank you

Answer by checkley79(3341)

4+5=9 WHICH REPRESENTS ONE GROUP OF STUDENTS.
1620/9=180 TOTAL GROUPS OF STUDENTS
THERE ARE 4 BOYS IN EACH GROUP.
4*180=720 TOTAL BOYS.
THERE ARE 5 GIRLS IN EACH GROUP.
5*180=900 TOTAL GIRLS.
PROOF:
720+900=1620
1620=1620
##############################################
HOPE THIS IS AN ADEQUATE SOLUTION METHOD.